### v2

# ECORE: Energy-Conscious Optimized Routing for Deep Learning Models at the Edge

Abstract.
Edge computing enables data processing closer to the source, significantly reducing latency—an essential requirement for real-time vision-based analytics such as object detection in surveillance and smart city environments. However, these tasks place substantial demands on resource-constrained edge devices, making the joint optimization of energy consumption and detection accuracy critical. To address this challenge, we propose ECORE , a framework that integrates multiple dynamic routing strategies—including estimation-based techniques and a greedy selection algorithm—to direct image processing requests to the most suitable edge device–model pair. ECORE dynamically balances energy efficiency and detection performance based on object characteristics. We evaluate our approach through extensive experiments on real-world datasets, comparing the proposed routers against widely used baseline techniques. The evaluation leverages established object detection models (YOLO, SSD, EfficientDet) and diverse edge platforms, including Jetson Orin Nano, Raspberry Pi 4 and 5, and TPU accelerators. Results demonstrate that our proposed context-aware routing strategies can reduce energy consumption and latency by 45% and 49%, respectively, while incurring only a 2% loss in detection accuracy compared to accuracy-centric methods.
Edge computing, Dynamic Routing, Object Detection Models, Energy Consumption, Accuracy.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: ; May 11-14, 2026; Saint-Malo, Franceisbn: XXX-X-XXXX-XXXX-X/2025/06
1. Introduction
The rapid growth of the Internet of Things (IoT) and smart applications in recent years has led to a massive increase in data generated by distributed sensing devices. According to a forecast by the International Data Corporation (IDC) (Reinsel, 2019), there will be 55.9 billion IoT devices generating approximately 79.4 zettabytes (ZB) of data by 2025. Continuously transmitting this volume of data to centralized data centers or cloud platforms for processing is often infeasible due to several limitations, including network bandwidth constraints, increased latency, and potential privacy and security concerns. These challenges are particularly significant in applications that require real-time decision-making or operate in bandwidth-limited environments. As a response, edge computing has emerged as a viable solution, enabling data processing to occur closer to the source–on or near the IoT devices using resource constrained single board computers such as Jetson, Raspberry Pi, and similar edge platforms. By offloading computation to the edge, this paradigm reduces communication overhead, lowers response times, and enhances the scalability and autonomy of intelligent systems.
Object detection, a key task in computer vision-based analytics, has seen widespread adoption across various domains including agriculture, sports analytics, transportation, healthcare, and smart cities. Recent advancements in machine learning and artificial intelligence (AI) have significantly improved the performance of object detection models. In particular, deep learning-based models such as You Only Look Once (YOLO) (Redmon et al., 2016), Single shot multibox detector (SSD) (Liu et al., 2016) and EfficientDet (Tan et al., 2020) have become increasingly prominent due to their high detection accuracy and real time capabilities. These models are often implemented using convolutional neural network (CNN) architectures and come in multiple variants to accommodate different computational budgets. For instance, the YOLOv8 (Jocher et al., 2023) family includes nano, small, medium, and large versions each reflecting trade-offs between model size, speed, and accuracy.
The dynamic nature of object detection in real-world environments presents both a challenge and an opportunity. Visual scenes vary widely in complexity—from sparsely populated streets at night to dense daytime urban intersections—leading to fluctuating computational demands for object detection. In such contexts, the heterogeneity of edge resources, including both hardware devices and detection models, offers an opportunity for intelligent allocation of requests to device–model pairs.
Our recent findings show that detection models exhibit notably different energy, latency, and accuracy profiles when deployed across various edge devices (Alqahtani et al., 2024). By leveraging this hardware–model heterogeneity, systems can achieve substantial energy savings without compromising detection accuracy or latency.
Energy efficiency is particularly critical on edge devices operating under strict power constraints, such as those relying on batteries or local energy harvesting methods like solar panels. Yet in safety-critical applications like surveillance and traffic monitoring, accuracy remains non-negotiable. This creates the central challenge addressed in this study: “how can we minimize energy usage without degrading detection quality?
The solution lies in adaptive, context-aware routing strategies that dynamically select the optimal device–model pair based on scene complexity. Simpler scenes allow for lightweight processing, while complex environments require more accurate–and consequently more energy-intensive–models. To address this, we propose the ECORE framework, which introduces a set of dynamic, novel, intelligent routing algorithms to redirect an image processing task to the most appropriate device-model pair. The goal is to reduce energy consumption while maintaining high detection accuracy.
ECORE’s routers operate by applying lightweight complexity estimation techniques during the pre-processing stage to infer the number of objects in each incoming image. Based on this estimation, the router selects and forwards the request to an optimal model–device pair.
The main contributions of this work are as follows:
  1. (1)
    We propose a complete system architecture for a framework called ECORE , optimized routing of image processing tasks to a group of heterogeneous edge devices running diverse object detection models, coordinated by a central gateway responsible for intelligent routing decisions.
  2. (2)
    We develop three object count estimation approaches to support intelligent routing decisions in ECORE . The first is an edge detection–based approach, which utilizes the Canny technique to estimate the number of objects in an image. The second is an SSD-based front-end approach, which deploys a lightweight object detection model at the gateway to make decisions. The third is an outputs-based approach, which leverages the object detection results returned from the backbend in the previous processing round and uses this information to inform routing decisions for the subsequent request.
  3. (3)
    We conduct a comprehensive experimental evaluation of our proposed routing approaches by comparing them against both an ideal benchmark–an Oracle Router that knows the exact number of objects in the image–and several baseline techniques, including Random, Round Robin, Lowest Energy Consumption, Lowest Inference Time, and Highest mean Average Precision (mAP). The evaluation is conducted on a real-world testbed using multiple datasets, including the Common Objects in Context (COCO) validation set and a short video, under diverse input scenarios.
The remainder of the paper is organized as follows: Section 2 discusses the motivation behind this work. Section 3 presents the system design, problem formulation, and greedy routing approach. Section 4 describes the datasets, experimental setup, and results. Section 5 reviews related literature. Finally, Section 6 concludes the paper and outlines directions for future work.
2. Motivation
This section highlights the motivation for developing adaptive, context-aware routing strategies in edge-based object detection systems. Figure LABEL:fig:scene_examples illustrates two contrasting scenes captured by an urban surveillance camera at an intersection. The first image depicts numerous pedestrians crossing the intersection, while the second image shows the same location when it is sparsely populated. In the former high density situation, accurate object detection becomes more critical and potentially more computationally demanding, especially in safety sensitive applications such as surveillance, traffic monitoring, and crowd management. Failure to detect or correctly classify objects in this context can compromise system reliability and lead to safety risks. In contrast, the latter image shows a sparsely populated street with minimal environmental activity and few objects of interests. In such low activity scenarios, the computational complexity of achieving reasonable accuracy is reduced, and the risk of critical misdetections is substantially lower. Therefore, energy savings can be achieved by using lighter detection models, particularly when edge devices operate under battery or solar power constraints.
These scenarios underscore the importance of a dynamic routing approach that forwards requests to the optimal edge node based on scene complexity. By estimating the complexity of a scene prior to routing, the system can intelligently choose between high accuracy models for crowded scenes and energy efficient models for sparse scenes. This motivation drives the design of our proposed routing algorithm, which aims to optimize energy consumption without compromising accuracy where it matters most.
To verify our hypothesis, we conducted a preliminary experiment using the COCO validation dataset, which contains 5,000 images. We categorized the images into two groups based on the number of objects, as determined by the ground truth: one group with a single object per image and another with four or more. We then calculated the mAP and energy consumption per image during inference. As shown in Figure LABEL:fig:energy_and_mAP(b), SSD Lite and YOLOv8n exhibit similar accuracy for single-object images, while YOLOv8n significantly outperforms SSD Lite on images with four or more objects—achieving nearly twice the mAP. In terms of energy usage, SSD Lite consumes a consistent amount across both groups, which is approximately 50% lower than YOLOv8n. These results suggest that in less complex scenarios, two different object detection models can achieve similar accuracy, while significant energy savings is possible by using more efficient models without compromising accuracy.
3. System Design and Implementation
This section describes the system architecture of ECORE , designed for energy-efficient and performance-aware routing of object detection tasks in an edge–IoT environment. Figure 2 illustrates the overall design. We assume a scenario where a pool of backend edge nodes hosts a variety of model–device pairs. Each backend device runs a dedicated object detection model and exposes an inference API to process images or video frames captured via distributed cameras or embedded sensors. This modular and distributed architecture ensures scalability and responsiveness, making it suitable for real-time applications in energy-constrained and performance-critical edge computing scenarios.
Cameras or embedded sensors transmit image data or video frames to a central gateway, which performs lightweight preprocessing and analysis. As shown in Figure 2, the gateway first applies one of the proposed object estimation techniques to infer the number of objects in each incoming image. This count is passed to the routing algorithm (Algorithm 1), which uses it to determine the most appropriate model–device pair.
The routing algorithm begins by identifying the corresponding object count group based on the estimated number of objects, using a predefined set of group rules defined by numeric ranges and labels. It then filters the profiling data to retain only the entries that match the identified group. From this subset, it calculates the maximum achievable mAP and derives a minimum acceptable mAP by subtracting a predefined margin. The data is then filtered again to retain only model–device pairs whose mAP falls within this target range. Among these, the pair with the lowest energy consumption is selected. The gateway forwards the request to the selected endpoint, which processes the image and returns the detected objects and their bounding boxes.
Figure 2. System architecture presentation.
3.1. Problem Formulation and Greedy Approach
When running an object detection application in an edge environment, a key challenge lies in routing image requests to the most suitable model-device pair, taking into account trade-offs among accuracy, latency, and energy consumption. Let
=$\mathcal{M}=\{(m_{1},d_{1}),(m_{2},d_{2}),\dots,(m_{n},d_{n})\}$
denote the set of available object detection model-device pairs, where $m_{i}$ represents a specific object detection model and $d_{i}$ denotes the device on which it can be executed. Each pair $(m_{i},d_{i})\in\mathcal{M}$ is profiled in advance, yielding key performance indicators:
  • $\text{mAP}_{i}$: mean Average Precision
  • $t_{i}$: inference time (in milliseconds)
  • $e_{i}$: energy consumption (in milliwatt-hour (mWh))
  • $g_{i}$: object count group (e.g. 0, 1, 2, 3, etc)
Given an incoming image request accompanied by an estimated object count group $G$, the routing objective selects the most energy-efficient model–device pair that satisfies application constraints for accuracy. This decision-making process can be formulated as a constrained optimization problem. First, the subset $\mathcal{M}_{G}\subseteq\mathcal{M}$ is identified, containing only the pairs belonging to group $G$. From this subset, the algorithm determines the maximum achievable accuracy.
$\text{mAP}_{\text{max}}=\max_{i\in\mathcal{M}_{G}}\text{mAP}_{i}$
To preserve flexibility in model-device selection and allow for more energy-efficient decision, we introduce a dynamic tolerance margin, $\delta_{\text{mAP}}$. This controlled relaxation enables the system to trade off a small reduction in mAP for gains in energy efficiency. The margin defines a feasible set $\mathcal{F}\subseteq\mathcal{M}_{G}$, within which model-device pairs are considered acceptable for routing, and formally defined as:
=$\mathcal{F}=\{i\in\mathcal{M}_{G}\mid\text{mAP}_{i}\geq\text{mAP}_{\text{max}}-\delta_{\text{mAP}}\}$
The routing algorithm then selects the optimal pair by minimizing energy consumption:
$i^{*}=\arg\min_{i\in\mathcal{F}}e_{i}.$
Algorithm 1 Routing Algorithm
0:  number_of_objects, profiling_data, $\delta$-mAP, group_rules
1:  Determine group by searching group_rules:
[…]
5:        break
6:     end if
7:  end for
[…]
9:  return group_data
10:  Compute max_mAP = maximum mAP (group_data)
[…]
13:  return refined_data
14:  Select the row in refined_data that has the lowest energy
15:  return the device–model
This greedy algorithm, as detailed in Algorithm 1, provides a principled balance between performance and efficiency, enabling the system to route requests to energy-efficient configurations with only marginal degradation in accuracy. It is particularly suitable for real-time applications deployed in resource constrained environments, where the router must make fast, intelligent decisions under uncertain and dynamic workloads.
3.2. Optimality of the Routing Algorithm
While the routing problem proposed in the previous section involves trade-offs among multiple objectives, applying threshold-based filtering prior to selection reduces the decision to a one-dimensional optimization. This simplification not only lowers computational complexity but also guarantees optimality under the model assumptions. As a result, the algorithm remains robust, fast, and well-suited for deployment in resource-constrained, real-time environments. The structure of the problem ensures that the greedy solution is indeed optimal. The following theorem formally proves the optimality of the greedy routing algorithm presented in Algorithm 1.
Theorem 3.1 (Optimality of the Routing Algorithm).
Let $\mathcal{M}_{G}$ be the subset of profiled model-device pairs matching the estimated object count group $G$, and let $\delta_{\text{mAP}}$ be a fixed accuracy tolerance. Then, the greedy selection of the model-device pair with the lowest energy consumption from the filtered feasible set
=$\mathcal{F}=\left\{i\in\mathcal{M}_{G}\mid\text{mAP}_{i}\geq\text{mAP}_{\max}-\delta_{\text{mAP}}\right\}$
yields an optimal solution to the routing problem under the specified constraints.
Proof.
The routing objective is to minimize energy consumption:
$\min_{i\in\mathcal{F}}e_{i}$
where the feasible set $\mathcal{F}$ includes only those model-device pairs in $\mathcal{M}_{G}$ whose accuracy exceeds the threshold $\text{mAP}_{\max}-\delta_{\text{mAP}}$.
Since all the constraints including accuracy and group relevance are applied during the filtering step, the remaining selection is unconstrained and reduces to choosing the element with minimum $e_{i}$.
As the energy consumption values $e_{i}$ are known from prior profiling and independent of each other, the greedy algorithm that selects
$i^{*}=\arg\min_{i\in\mathcal{F}}e_{i}$
is guaranteed to return the globally optimal solution to the constrained minimization problem. There are no interdependencies or temporal dynamics that would invalidate this selection.
3.3. Proposed object count estimation approaches
It is crucial to keep the core of the routing algorithm proposed in the previous subsection—particularly the component responsible for estimating the number of objects in an image—lightweight. Otherwise, the overhead introduced by the algorithm could offset the energy savings achieved by the system. Therefore, to estimate the number of objects in incoming image requests and redirect them to the most suitable edge node, we develop and evaluate three estimation approaches:
  1. (1)
    Edge Detection (ED): This approach leverages edge detection which is a fundamental technique in image processing and computer vision used to identify areas of significant intensity or color change, which often correspond to object boundaries. The primary goal is to capture essential structural information. We employ the widely used Canny edge detection algorithm (Canny, 1986), to estimate the number of objects in each image. The resulting object count is then fed into the proposed routing algorithm (Algorithm 1), which selects the optimal model-device pair based on the expected energy–accuracy trade-off governed by $\delta_{\text{mAP}}$. This method is considered lightweight and computationally efficient, though it may yield coarse or approximate estimations.
  2. (2)
    SSD-Based Front-End (SF): In this approach, we deploy a very lightweight and efficient Single Shot MultiBox Detector (SSD) object detection model directly at the gateway. The model processes incoming images to identify and count objects, and the resulting count is passed to the routing algorithm, which determines the appropriate edge node. Compared to the edge detection method, this approach is expected to provide more accurate object count estimates, albeit at a higher computational cost at the gateway.
  3. (3)
    Output-Based (OB): This approach leverages temporal continuity in image streams—particularly video—by assuming that consecutive frames contain a similar number of objects. Instead of estimating the object count for every image (frame), the system reuses the count obtained from the output of the previous image (frame), as detected by the last selected model–device pair. The same configuration is retained until a change in object count is observed. The routing algorithm begins with a default estimate (e.g., zero) and forwards the first image to an edge node. Once the detection results are returned, the router extracts the actual number of objects and uses it for subsequent routing decisions. If the object count remains stable, the system continues using the same model–device pair; if a change is detected, the routing decision is updated accordingly. This adaptive strategy avoids per-image (frame) estimation, reducing computational overhead and improving energy efficiency. OB is particularly effective for video streams or sequences of images that exhibit consistent object counts or computational complexity, where the number of objects tends to vary gradually over time.
4. Performance Evaluation
4.1. Experimental Setup
4.1.1. Dataset
In this study, we utilize three distinct datasets, two of which are derived from the COCO validation set (Lin et al., 2014), to evaluate the routing algorithm. The first is the Original COCO Validation Dataset, consisting of 5,000 images annotated with ground truth labels, including object classes and bounding boxes spanning 80 object categories. As this dataset reflects natural scene variability, the number of objects per image varies significantly. Figure 3 illustrates the distribution of object counts, highlighting this diversity.
The OB routing algorithm is designed to prioritize continuity, favoring sequences of visually similar images with minimal variation in object density. To support this principle, we constructed a secondary dataset from the COCO validation set, termed the Balanced Sorted Dataset. This dataset consists of 1,000 images, evenly distributed across five categories based on the number of objects per image: ‘0’, ‘1’, ‘2’, ‘3’, and ‘4 or more’. Each category contains 200 images selected to maintain uniform object counts within groups. Specifically, Group 1 comprises images with no objects; Group 2, exactly one object; Group 3, two objects; Group 4, three objects; and Group 5, four or more objects. When a group contains fewer than 200 unique images, random duplications are introduced to reach the target count. The images are then sent to the gateway sequentially, ordered by group.
We also evaluated our routing pipeline on a real-world video stream of pedestrians crossing the road by treating it as a detection dataset. Because no manual annotations were available, we decoded the raw file and extracted every frame (at the original frame rate) into a sequence of JPEG images. We then generated the ground-truth annotations by running the powerful YOLOv8x model over each frame—saving its detections in YOLO-format text files to serve as our reference labels.
Figure 3. Distribution of Object Counts per Image in the COCO Validation Set.
4.1.2. Testbed
This section outlines the experimental testbed used to evaluate the proposed routing approach. The testbed configuration is informed by extensive prior benchmarking results (Alqahtani et al., 2024), which profiled various combinations of edge devices and object detection models across three key performance metrics: energy consumption, inference time, and mAP. The devices evaluated included Raspberry Pi 3 (with and without TPU), Raspberry Pi 4 (with and without TPU), Raspberry Pi 5 (with and without TPU and with an AI HAT), and the Jetson Orin Nano, making a total of eight. Eight object detection models profiled ranged from lightweight variants—SSD v1, SSD Lite, EfficientDet-Lite0, Lite1, and Lite2 to the YOLOv8 family (Nano, Small, and Medium). These metrics collectively reflect the trade-offs inherent in deploying object detection models on edge hardware. Notably, the benchmarking revealed that no single model–device pair consistently outperforms others across all criteria. Therefore, model selection must be context-dependent, guided by the specific performance objective. Figure 4 presents the accuracy–energy consumption trade-offs for all possible 64 model–device combinations.
To construct the testbed, we selected only model–device pairs that lie on or near the Pareto front in at least one performance dimension. We repeated the all the configurations in Figure 4 per 5 object groups, i.e., ‘0’, ‘1’, ‘2’, ‘3’, and ‘4 or more’ objects per image. This ensures that all included devices offer meaningful trade-offs—i.e., each is strong in at least one metric—while excluding configurations that are consistently suboptimal across all criteria. Energy efficiency and latency do not vary across groups. Therefore, for energy efficiency, we selected Jetson Orin Nano running SSD v1 which demonstrated the lowest energy consumption among all configurations. In terms of inference latency, the Raspberry Pi 5 with TPU paired with SSD v1 achieved the shortest inference time. For detection accuracy (mAP), the optimal configuration varied across object count groups, underscoring the need for adaptive routing. Specifically, for Group 1 (images with no objects), the Raspberry Pi 5 with TPU running SSD v1 achieved the highest mAP. For Group 2 (images with one object), the Raspberry Pi 5 with TPU running SSD Lite performed best. In Group 3 (images with two objects), the Jetson Orin Nano with YOLOv8-small yielded the highest accuracy. Finally, for Groups 4 and 5 (images with three and four or more objects), the Raspberry Pi 5 equipped with an AI Hat running YOLOv8-small demonstrated superior performance. Table 1 provides the finalized list of device-model pairs, and Table 2 details the corresponding device specifications. This testbed supports a thorough assessment of the routing algorithm’s effectiveness in achieving optimal trade-offs between energy efficiency and performance across a variety of edge computing environments.
Figure 4. Pareto frontiers comparing object detection models across various edge devices.
Table 1. Experimental Testbed Configurations
Metrics Edge Device Object Detection Model Framework
Energy Consumption Jetson Orin Nano SSD v1 TensorRT
Inference Time Raspberry Pi 5 SSD v1 TFLite
mAP – Group 1 Raspberry Pi 5 + TPU SSD v1 TFLite
mAP – Group 2 Raspberry Pi 5 SSD Lite TFLite
mAP – Group 3 Jetson Orin Nano YOLOv8-small TensorRT
mAP – Group 4 Raspberry Pi 5 + AI Hat YOLOv8-small HEF
mAP – Group 5 Raspberry Pi 5 + AI Hat YOLOv8-small HEF
Table 2. Testbed Device Specifications
Device Name Processor Memory OS/SDK Bluetooth Wi-Fi
Raspberry Pi 4 CPU 4 GB Debian Bookworm Off Off
Raspberry Pi 5 CPU 4 GB Debian Bookworm Off Off
Pi 5 + Coral TPU TPU 4 GB Debian Bookworm Off Off
Pi 5 + AI Hat Hailo-8 4 GB Debian Bookworm Off Off
Jetson Orin Nano GPU 8 GB JetPack 5.1.3 Off Off
4.2. Experimental Procedure
To the best of our knowledge, no existing work in the literature directly addresses the same problem with a comparable approach. Therefore, to rigorously evaluate the effectiveness of our proposed routing strategies, we systematically compared them against a set of baseline routing methods, including an idealized benchmark referred to as the Oracle Router, as well as several widely used strategies: Round Robin, Random, Lowest Energy, Lowest Inference Time, Highest mAP an Highest mAP per Group. The baseline and benchmark routing strategies are defined as follows:
  • Oracle Router (Orc): Serves as an ideal benchmark by assuming perfect prior knowledge of the number of objects in each incoming image. This is achieved by leveraging the ground truth annotations. The true object count is passed as metadata alongside the image request to the gateway, enabling the routing algorithm to select the optimal model–device pair by applying the $\delta$mAP criterion. This router reflects the upper bound of routing performance under perfect estimation.
  • Round Robin Router (RR): Distributes incoming image requests sequentially in a round robin fashion across all available nodes, regardless of their energy efficiency or performance.
  • Random Router (Rnd): Assigns each request to a node randomly, independent of any system metrics or workload history.
  • Lowest Energy Router (LE): Routes each image to the model–device pair that consumes the least energy, prioritizing power efficiency over accuracy or latency.
  • Lowest Inference Time Router (LI): Selects the node with the shortest average inference time, aiming to minimize latency without considering energy or accuracy.
  • Highest mAP Router (HM): Always chooses the model–device pair with the highest mAP, independent of the object count in the image.
  • Highest mAP per Group Router (HMG): Within each object-count group, the model–device pair achieving the highest mAP is selected, irrespective of energy consumption or latency constraints.
Given the varying performance requirements of different real-world applications, we evaluated the proposed and baseline routers across a range of $\delta$mAP values (e.g., 0, 5, 10, 15, 20, and 25). A value of $\delta$mAP = 0 enforces strict accuracy, whereas higher values allow greater flexibility in selecting less accurate but more energy-efficient models.
To ensure consistent load generation, we used Locust (Jonatan et al., 2025) to send image requests back-to-back in a piggybacked fashion—each new request is triggered only after the previous image has been processed and a response is received, simulating continuous workloads. Performance was evaluated using the following primary metrics:
  • Energy Consumption: Measured in milliwatt-hours (mWh) over the full duration of the experiment. To isolate the energy used for processing, we subtracted the baseline idle energy consumption of all devices when no tasks are executed but the devices remain powered on.
  • Latency: Reported in seconds as the total time required to complete all image requests from start to finish.
  • Accuracy: Evaluated using the mAP, computed via the FiftyOne tool (Voxel51, 2025), which provides robust comparisons against ground truth annotations.
  • Gateway Overhead: To assess the computational cost and overhead of different routing strategies, we separately report the energy consumption (mWh) and latency (seconds) incurred within the gateway during the routing decision-making process. This metric isolates the overhead introduced by each router, highlighting the trade-offs between routing complexity and performance.
Experiments were conducted using the three datasets described earlier: (1) the original COCO validation set, which naturally reflects varied object distributions; (2) the balanced and sorted dataset was constructed to include an equal number of images for each object count group, with images grouped by object count—i.e., all images containing zero object first, followed by those with one object, and so on—thus collectively favoring the output-based approach; and (3) a pedestrian-crossing video dataset, where every frame of a real-world road-crossing sequence was extracted as an image. This triple-dataset evaluation enables a comprehensive analysis of routing behaviour under both real-world and controlled conditions.
4.3. Experimental Results
4.3.1. Full COCO Dataset Results:
To evaluate the performance of the proposed routing approaches in comparison to baseline methods, we conducted experiments using the full COCO validation dataset, with the $\delta$ mAP threshold set to 5 unless otherwise specified. The results across energy consumption, latency, and detection accuracy are presented in Figure LABEL:fig:results_full_coco_dataset. In terms of dynamic energy consumption Figure LABEL:fig:results_full_coco_dataset(b), LE achieves the lowest energy usage, establishing the theoretical lower bound. LI follows closely, consuming only marginally more. On the other end of the spectrum, HMG, achieving best accuracy per group, increases energy consumption by approximately 80 percent over the lower bound. The proposed SF exhibits an even greater rise of nearly 85%, attributed to the additional overhead introduced by gateway analysis. The Orc router results in only 50% increase in energy usage. Among the proposed methods, ED and OB increase consumption by about 45% and 37%, respectively. In comparison, RR and Rnd lead to energy increases of around 55% and 50%.
With respect to detection accuracy (mAP), Figure LABEL:fig:results_full_coco_dataset(a) shows that HMG achieves the highest score, setting the upper bound. Orc and SF perform similarly, with accuracy dropping by less than one percent. HM and ED experience slight reductions of approximately 1% and 2%, respectively. A more significant 9% drop is observed in the OB router’s performance, attributed to the highly variable and unpredictable object counts present in the COCO dataset. Finally, RR and Rnd incur significant drops of around 25%, while LE and LI show the largest losses—about 40% and 50%, respectively. Results show that the proposed routing approaches, such as ED and OB, can achieve mAP values very close to optimal methods like HMG, while also delivering significant energy savings of approximately 45% and 37%, respectively.
Latency results, measured as the total time to process the entire dataset, are presented in Figure LABEL:fig:results_full_coco_dataset(c). As expected, LI achieves the lowest overall latency, with LE following closely—incurring an increase of less than 1%. In contrast, HM and SF exhibit the highest latency increases, each rising by approximately 75%, while HMG incurs a latency overhead of around 60%.. In contrast, Orc, OB, and ED introduce relatively small latency increases of about 11%.
Insight #1. Considering the upper bound on achievable mAP (as given by HMG) and the minimum energy consumption (as achieved by LE), the proposed ED approach offers the most efficient trade-off on static-image datasets such as COCO. It attains accuracy close to the oracle-level performance—which has full knowledge of object counts—while maintaining competitive energy efficiency and incurring only minimal latency overhead.
4.3.2. Balanced Sorted Dataset Results
This section highlights the results of evaluating our proposed routing approaches and baseline routing strategies using the balanced sorted dataset as shown in Figure LABEL:fig:results_sorted_balanced_dataset. With respect to dynamic energy consumption Figure LABEL:fig:results_sorted_balanced_dataset(b), LE demonstrates the most energy-efficient performance, consuming only 227 mWh and establishing the theoretical lower bound. LI performs similarly, with only a slight increase in consumption. At the opposite end, HMG exhibits a substantial rise in energy usage—approximately 50% higher than the baseline—while SF records more than double the energy consumption of LE. The Orc router shows a 40% increase in energy use relative to the lower bound. Among the proposed approaches, ED and OB consume 64% and 45% more energy, respectively. In comparison, the baseline methods RR and Rnd result in increases of roughly 70% and 62%.
Latency results, presented in Figure LABEL:fig:results_sorted_balanced_dataset(c), show that LI achieves the lowest total latency at approximately 306 seconds, with LE following closely behind, incurring only a 2-second increase. HM and SF exhibit the highest latency overheads, with increases of around 77% and 81%, respectively. HMG also shows a notable rise, adding approximately 34% to the baseline latency. In contrast, Orc and OB introduce modest increases of about 9%, while ED shows a slightly higher increase of 13%. The RR and Rnd routers fall in the mid-range, with latency increases of 18% and 22%, respectively.
In terms of detection accuracy (mAP), Figure LABEL:fig:results_sorted_balanced_dataset(a) indicates that HMG achieves the highest score at 40.94, establishing the upper performance bound. Orc, SF, and OB closely follow, each exhibiting a negligible decline of less than 1%. ED also shows minor reductions of approximately 1%. In contrast, RR and Rnd suffer notable accuracy losses of around 18%, while LE and LI demonstrate the most substantial declines, with reductions of approximately 30% and 40%, respectively.
Insight #2. Among the proposed routing methods, OB offers the best overall trade-off, closely approaching oracle-level performance in both accuracy and latency while achieving the lowest energy consumption of all proposed approaches. This makes OB the most suitable choice for sequential image processing scenarios.
4.3.3. Video Dataset Results
This section presents the results of the different routing approaches evaluated using the pedestrian video dataset. The corresponding Figure LABEL:fig:results_video_dataset illustrates the performance trends across key metrics. With respect to dynamic energy consumption Figure LABEL:fig:results_video_dataset(b), LE consumes only 85 mWh. LI follows with a 75% increase in consumption over LE. At the opposite end, HMG shows a substantial rise in energy usage—approximately 100% higher than the baseline—while SF consumes more than three times the energy of LE, exceeding 200%. The Orc router results in an 85% increase relative to the lower bound. Among the proposed methods, ED and OB increase energy usage by approximately 138% and 100%, respectively. For comparison, the baseline methods RR and Rnd show similar trends, each exceeding 200% of the energy consumption of LE.
Latency results, illustrated in Figure LABEL:fig:results_video_dataset(c), indicate that LI achieves the best performance, with the lowest total latency of approximately 164 seconds. LE follows closely, incurring a minimal increase of 7 seconds. On the higher end, HM and SF experience the greatest latency overheads, with increases of approximately 61% and 73%, respectively. HMG also shows a significant rise, adding around 50% over the baseline. In comparison, Orc and OB introduce only modest latency increases of about 9%, while ED results in a more pronounced increase of 40%. The RR and Rnd routers occupy a middle ground, with latency overheads of 29% and 17%, respectively.
Regrading accuracy (mAP), Figure LABEL:fig:results_video_dataset(a) shows that HMG achieves the highest score at 79.95, setting the upper bound for performance. Orc closely follows, with a minimal decline of less than 1%. SF, OB, and HM each experience moderate reductions of around 4%. ED shows a more significant drop, with accuracy decreasing by approximately 14%. In contrast, RR and Rnd incur substantial losses of 50% and 55%, respectively, while LE and LI show the most severe accuracy degradation, with mAP reductions of about 63% and 75%.
Insight #3. Among the proposed methods, OB emerges as the most effective for video scenarios, combining high accuracy with low latency and reasonable energy consumption—benefiting from the continuity and temporal consistency of video frames.
4.3.4. Results Across Different $\delta$mAP Values
This section presents the performance of the Oracle and proposed routers across varying $\delta$mAP values. All routers including the Oracle show a substantial reduction in energy consumption when $\delta$mAP increases from 0 to 5. Between $\delta$mAP values of 5 and 20, energy consumption continues to decrease slightly, followed by a more noticeable drop at $\delta=25$ as Figure LABEL:fig:results_differnt_delta depicts.
A similar trend is observed for total latency. All routers exhibit a sharp decline in latency when $\delta$mAP increases from 0 to 5, with only marginal improvements observed between values of 5 and 25.
In terms of detection accuracy (mAP), most routers maintain stable performance as $\delta$mAP increases from 0 to 5. However, a slight decrease in mAP is observed between $\delta=5$ and $\delta=15$, followed by a more significant decline between $\delta=20$ and $\delta=25$.
Insight #4. Results show that even setting $\delta$mAP to a small value such as 5—allowing up to a 5% tolerance in accuracy drop—yields significant reductions in both energy consumption and latency, while maintaining detection performance comparable to the strictest setting ($\delta=0$). Notably, with $\delta$mAP set to 5, the actual accuracy drop remains relatively low, around 2%.
4.4. Discussions
ECORE demonstrates promising results by reducing energy consumption and latency while maintaining strong accuracy. While the Orc router represents an idealized baseline, benefiting from prior knowledge of object counts, achieving comparably strong results in practice is not infeasible. Our proposed routing approaches—ED and OB—achieve performance close to Oracle, with marginal differences. However, the effectiveness of each method is dataset-dependent: OB performs optimally in sequential and structured scenarios, such as the sorted balanced and video datasets, while ED is better suited for static image inputs without any continuity.
The current design assumes static profiling, where performance metrics remain stable at runtime, which may not reflect real-world dynamics such as temperature, background load, or battery state. Additionally, our heuristic operates at single-request granularity, limiting its applicability in batch or load-balancing contexts. It is important to note that the proposed approaches do not support multi-objective optimization; balancing trade-offs between energy and latency would require a Pareto-optimal or weighted strategy, where a greedy algorithm may no longer suffice.
5. Related Work
Table 3. Comparison of Related Work and ECORE
Study Edge Com. Obj. Det. Dynamic Routing Energy Cons. Accuracy Real Testbed
Ji et al. (Ji et al., 2025)
Trinh et al. (Trinh et al., 2018)
Tu et al. (Tu et al., 2023)
Zhang et al. (Zhang et al., 2022)
Tundo et al. (Tundo et al., 2023)
Matathammal et al. (Matathammal et al., 2025)
Kulkarni et al. (Kulkarni et al., 2023)
Marda et al. (Marda et al., 2024)
Stripelis et al. (Stripelis et al., 2024)
Maurya et al. (Maurya et al., 2024)
Zheng et al. (Zheng et al., 2025)
Guha et al. (Guha et al., 2024)
Mohammadshahi et al. (Mohammadshahi et al., 2024)
Sikeridis et al. (Sikeridis et al., 2024)
ECORE
This section discusses related studies and compares them with our proposed work as Table 3 depicts. Ji et al. (Ji et al., 2025) introduce a Cloud-Edge Large-Tiny Model Collaborative (CELTC) architecture that employs deep reinforcement learning to enhance surveillance data processing and decision-making in offshore drilling platform environments. The CELTC framework integrates edge–cloud computing by deploying high-accuracy large models on cloud servers and lightweight, efficient models on edge devices to balance computational demand and responsiveness. Similarly, Trinh et al. (Trinh et al., 2018) explore the use of Mobile Edge Computing (MEC) to address energy management challenges in power-constrained IoT devices while enabling low-latency processing of high-resolution visual data. Focusing on facial recognition for disaster response scenarios, the authors propose an offload decision-making algorithm that dynamically determines whether to process tasks at the edge or offload them to the cloud, based on current workload and energy constraints. While these studies present valuable edge–cloud collaborative architectures, their focus differs from ours. Specifically, they rely on cloud resources to support edge processing, whereas our work emphasizes fully distributed collaboration among multiple edge devices. Furthermore, our approach directly targets the optimization of energy consumption and latency in edge environments without compromising detection accuracy, making it well-suited for real-time applications where both efficiency and performance are critical.
Tu et al. (Tu et al., 2023) propose EdgeLeague, a framework designed to manage multiple video streams with diverse Quality of Service (QoS) requirements, ensuring reliable surveillance performance under constrained edge resources and fluctuating uplink bandwidth. EdgeLeague leverages collaborative edge computing and optimizes camera network configurations to solve an NP-hard integer nonlinear problem by dynamically adjusting camera resolutions and selecting appropriate detection models across cooperating edge nodes. Similarly, Zhang et al. (Zhang et al., 2022) address configuration selection and bandwidth allocation for multiple video streams connected to a shared uplink channel through a common edge node. The authors propose JCAB, an efficient online algorithm that jointly optimizes configuration adaptation and bandwidth distribution. The approach is further extended to a multi-edge scenario, allowing individual users or video streams to select their preferred edge node, thereby enhancing flexibility and resource utilization. While these studies focus on optimizing system performance by adapting hardware configurations and bandwidth allocation, our approach takes a different direction. Instead of adjusting device settings, we analyze the content of incoming images and dynamically route requests to the most suitable edge device–model pair. This enables energy-efficient inference while preserving detection accuracy, offering a lightweight and scalable alternative to hardware-level reconfiguration strategies.
Tundo et al. (Tundo et al., 2023) introduces an energy-conscious strategy for designing and deploying self-adaptive, AI-driven applications, aiming to balance application-level objectives—such as object detection accuracy and frame processing rate—with energy consumption. The authors address the challenge of identifying optimal system configurations for self-adaptation using a meta-heuristic search method that requires minimal empirical sampling. Similarly, Matathammal et al. (Matathammal et al., 2025) present EdgeMLBalancer, a self-adaptive methodology designed to optimize CPU utilization and resource management on edge devices. Their approach dynamically switches between deployed models based on real-time CPU usage monitoring across processor cores. In Kulkarni et al. (Kulkarni et al., 2023), the concept of a Machine Learning Model Balancer is proposed to manage uncertainty in ML model outputs through the coordinated use of multiple models. Building on this idea, the authors introduce AdaMLS, which enhances the traditional MAPE-K (Monitor-Analyze-Plan-Execute over a Knowledge base) loop to enable continuous adaptation of machine learning systems. The study in Marda et al. (Marda et al., 2024) introduces SWITCH, a framework designed to improve system-level adaptability by enabling runtime model switching. SWITCH functions as a general-purpose web service applicable to various ML scenarios, and is specifically demonstrated in an object detection use case. Whereas these studies focus on self-adaptive techniques that switch between models within a single edge device, our work adopts a broader architectural perspective. We design a system that maintains a pool of edge devices and model variants, and introduces a dynamic routing mechanism to select the most appropriate device–model pair for each image. This approach supports distributed, energy-efficient processing across multiple edge nodes while preserving detection accuracy, extending adaptability beyond local device-level switching to collaborative edge inference.
Recent research has explored various routing strategies to improve the efficiency and adaptability of Large Language Model (LLM) inference. Stripelis et al. (Stripelis et al., 2024) introduce a modular querying system that dynamically dispatches incoming queries to specialized LLM experts based on the query’s requirements. Similarly, Maurya et al. (Maurya et al., 2024) propose an algorithm that selects the most appropriate subset of LLMs from a larger model pool to optimize performance and cost. Zheng et al. (Zheng et al., 2025) present CITER, a fine-grained token-level routing framework that enables collaborative inference between small and large language models by directing low-importance tokens to smaller models for efficiency and reserving complex tokens for larger models to maintain accuracy. Guha et al. (Guha et al., 2024) introduce Smoothie, a routing method based on weak supervision that constructs a latent variable graphical model from the embeddings of multiple LLM responses, enabling accurate output estimation without labeled data. Mohammadshahi et al. (Mohammadshahi et al., 2024) offers Routoo, a flexible routing architecture that balances latency, cost, and inference quality by allowing users to control trade-offs according to task requirements. Sikeridis et al. (Sikeridis et al., 2024) propose PickLLM which employs a reinforcement learning–based framework that dynamically assigns queries to models using a reward function guided by query-specific metrics such as accuracy, latency, and cost. While these works demonstrate significant advances in LLM routing, they are primarily focused on natural language processing tasks in cloud-based environments. In contrast, our work addresses object detection in edge computing settings, with a particular emphasis on minimizing energy consumption and processing latency and accuracy factors that are critical for real-time, resource-constrained applications.
6. Conclusions and Future Direction
Deploying object detection models on resource-constrained edge devices—particularly those powered by batteries or sustainable energy sources such as solar panels—requires careful consideration of energy consumption. In many real-world applications, especially safety-critical systems, the operating environment is highly dynamic, and the number of objects captured in each image can vary significantly. This variability can substantially affect the accuracy of object detection models, necessitating adaptive strategies that balance efficiency and performance.
To address these challenges, we proposed ECORE along with three dynamic, estimation-based routing algorithms that first estimate the number of objects in an image using lightweight pre-processing techniques. Then, based on this estimation, the algorithms route the request to the most suitable edge model-device pair. The routing decision is informed by profiling and is designed to minimize overall energy consumption while preserving detection accuracy within acceptable range.
Our experimental evaluation demonstrates that the proposed routing strategy performs favorably compared to baseline approaches. Specifically, it achieves a substantial reduction in energy usage while maintaining accuracy levels within an acceptable margin. The results highlight the potential of dynamic, context-aware routing for improving the energy-efficiency of AI workloads at the edge.
Future Work: To enhance the robustness of ECORE , future work will explore dynamic profiling to account for runtime variability such as temperature, battery state, and background load. We also aim to extend the routing strategy to support batch-level decision-making for better load balancing. Additionally, incorporating multi-objective optimization techniques, such as Pareto-based or weighted approaches, will allow more flexible trade-offs between energy consumption and latency across diverse deployment scenarios.